OpenStudy (anonymous):

Find the solution of y = 6x + 1 for x = 5

OpenStudy (anonymous):

y = 6.5 + 1 = 30 + 1 = 31

OpenStudy (slaaibak):

substitute x=5 in the equation.

OpenStudy (anonymous):

A. (5, 2/3) B. (–29, 5) C. (5, 36) D. (5, 31)

OpenStudy (anonymous):

y = 6x + 1 y = 6(5) + 1 y = 30 + 1 y = 31 y = 6x + 1 31 = 6x + 1 30 = 6x 30/6 = x 5 = x answer : (5,31)

OpenStudy (anonymous):

get off kelliegirl3

OpenStudy (anonymous):

huh ? What did I do ?

OpenStudy (anonymous):

I was just trying to help...sorry for whatever I did

OpenStudy (anonymous):

i just ask for ur # so we can text that's all

OpenStudy (anonymous):

A. y = -4x + 13 B. y = -4x – 13 C. y = 4x – 13 D. y = 1/4x – 13

OpenStudy (anonymous):

I am sorry, but I can't......I have a boyfriend and I don't want to cause any problems